Project

General

Profile

NDOS Commissioning Meeting 2011/09/29

Attendance: Jon, Peter, Mat, Gavin, Mark Mathis, Enrique, Gary, Tom, Susan, Jan, Minerba, Zeb, Rafael, Denis, Mark, Athans

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & Detector Update

  • On Tuesday (9/27/11), Walter Jaskierny will be removing his power monitoring equipment from rack 2 in the NDSB. This will require powering down that rack, thus the HV system. Jon will try to arrange for this to happen at 9 am. Martin: Gennadiy should be available in case there are problems. Jon will follow up with Peter
  • Mat: removed 15 APDs from 1st block and moved them to the verticals modules on the 2nd block. Also installed the "mechanical" coated APD on 1-2-31. This is operating warm for now; once we've verified that it's working correctly, we will attempt to cool it.
  • Mark: ~10 people have responded with a strong preference for "config. 4"
  • Jon: we will operate the recently-moved APDs for ~1 week. So unless there are major problems, we will install the next batch beginning on Monday Oct. 10
  • Mat: new installation technique went well. Torque-driver worked well, as well as the boroscope to get a better view of the APD installation. Problems were in fact found using the boroscope, so this was quite effective.
  • Mat: how will we monitor the APDs that were just moved? Jon: will ask shifters to do a DSO scan once during their shift.

DCS Update

DAQ Update

AOB

  • Jarek's Nearline Checklist is now available for shifters. Jon will add this to the "What to do while on shift" wiki.